# Verifying LLM Inference to Detect Model Weight Exfiltration

### v2

@@ 4 Theoretical Framework: Verification System
Figure 3: Overview of the verification system for runtime output validation. A lock symbol denotes parts of the inference verification scheme that require high integrity, further explored in Section D
@@
    Inference Logging. The API Server logs all inference inputs and outputs, together with references to the exact model release, code version, and the seed used for that inference.  [^4] In the event that not every (input, output, seed) tuple is verified, randomness is critical in the process of selecting which messages to verify is essential. If the messages-to-be-verified is predictable, an adversary could selectively encode sensitive data in unverified outputs..
@@
At the core of this architecture are three subsystems:
  1. 1.
    Trusted Inference Logging: Captures all the input message, output message, and seed tuples for verification.
  2. 2.
    Message Sampling Seed Generation: When not verifying every single output message, there needs to be a trusted scheme to generate randomness for selecting which inference messages to verify, which is neither predictable nor manipulable by an adversary.
  3. 3.
    Verification: Executes replays in a trusted, isolated environment to confirm distributional consistency with the reference model.
The degree of isolation of the Verification Server is a tunable security parameter. In minimal deployments, this can be enforced through access controls and containerization; in the extreme, the verification server may be fully air-gapped. All trust assumptions are detailed in Appendix D.
4.1 Message Sampling Rates
To understand the efficiency of our verification scheme, we analyze what percentage of traffic must be sampled to detect malicious messages with high confidence. Figure 4 illustrates the key statistical property of our approach: the sampling percentage required depends primarily on the proportion of malicious messages in the traffic. In Figure 4 we show the percentage of messages that need to be sampled in order to sample at least 1000 malicious messages, assuming the malicious messages are IID. Notably, even for very slow exfiltration attempts (e.g., 1% malicious traffic), the sampling percentage remains practical, even for high confidence. See Appendix C for the full derivation.
4.2 Solution 1: Perplexity and Syntactic Filters
One simple attack idea would likely be to exfiltrate weights directly through output messages, by sending the model weights as plaintext. A first line of defense against such a naive weight exfiltration technique is to detect messages that deviate from ordinary conversational patterns. Two simple, independent approaches can be implemented here.
Syntactic filtering. A basic scheme applies regular expressions and lightweight heuristics to flag messages with suspicious surface patterns—such as long numeric runs, base64 strings, or other structured encodings. This approach is trivial to deploy and computationally inexpensive, but is easily evaded by re-encoding information (e.g., mapping numbers to words or distributing bits across multiple tokens).
Perplexity filtering. A slightly more sophisticated approach uses a reference language model trained on benign message traffic to compute sequence-level perplexity—a standard measure of how well a model predicts a sequence. Formally, perplexity is the exponential of the average negative log-likelihood assigned to each token; lower values indicate that the model finds the text highly predictable, while higher values suggest the text is surprising or out-of-distribution. Messages with unusually high or low perplexity—indicating atypical token statistics—are flagged as potential exfiltration attempts [perplexityfilter]. This can detect outputs that are syntactically normal but statistically implausible under the benign distribution, though adaptive adversaries can still craft messages that maintain natural perplexity while encoding information.
Both schemes are simple to run and provide a low-cost first defense, but they leave a substantial gap between benign and adversarial outputs. Closing this gap requires defenses that consider semantics, context, or multi-turn correlations rather than surface-level or statistical anomalies alone.
4.3 Solution 2: Sampled-Token Verification via Exact Matching
@@
For efficiency, the same idea can be executed without iterative decoding. Using prefill, a single forward pass over the input+output yields the next-token distributions at every position; sampling (with the recorded seed) and exact matching can then be done position-wise in parallel. This preserves teacher-forcing semantics while running significantly faster. Full pseudocode appears in Appendix B.3.
While Inverse-Probability Transform and Gumbel-Max Trick are mathematically deterministic, in practice the logit distribution $\vec{p}_{\theta}(x)$ is produced by a non-deterministic computation. Even with a fixed seed $\sigma$, floating-point imprecision, asynchronous GPU execution, and nondeterministic kernel scheduling can cause small but measurable differences in $\vec{p}_{\theta}(x)$ across runs. This effect arises even in benign settings, independent of any adversarial manipulation, and has been well-documented in prior work [yuan2025fp32deathchallengessolutions]. We refer to this inherent variability as valid non-determinism.
$\vec{p}_{\theta}(x)\sim\mathcal{P}_{\theta}(x),$
where $\mathcal{P}_{\theta}(x)$ represents the stochastic process generating the token-level probabilities. This perspective allows us to define heuristic distance measures between two runs of the same model (with identical seed $\sigma$) by comparing samples from $\mathcal{P}_{\theta}(x)$—illustrated conceptually in Figure 5.
Fixed-Seed Posterior Distribution.
$t=\mathrm{Sample}(p,\sigma).$
This represents the probability—over draws of the underlying probability distribution pp—that a fixed random seed $\sigma$ yields token $t$. Intuitively, it characterizes how much stochastic variability remains in the sampling outcome even when the seed is fixed, due to non-determinism in the generation of $p$.
$P_{i}(t)=\Pr[T=t\mid X=x_{i}].$
$S_{i}(\sigma)=\min\{\,t\in\mathcal{V}:F_{i}(t)\geq R(\sigma)\,\},$
Intuitively, this distribution captures how much apparent randomness remains even when the seed is fixed. Figures 5(a) and 5(b) visualize the resulting fixed-seed posterior distributions for the inverse probability transform and the Gumbel-Max mechanism, respectively. These distributions form the basis for reasoning about an honest divergence during verification (as opposed to potentially malicious nondeterminism).
Understanding the limits of valid non-determinism is central to limiting the exfiltration rate of adversaries. Under very minimal trust assumptions (Appendix D), the only communication channel open to the adversary is by hiding communication within this honest variability.
4.5 Solution 3: Classification using the Fixed-Seed Sampling Likelihood Policy
The preceding section introduced the fixed-seed posterior distribution, which captures the stochastic variability of model outputs that can arise under honest non-determinism. Building on this, we define a practical decision rule for distinguishing between valid and suspicious generations. Our approach centers on estimating the likelihood that a given output token $t$ was honestly generated under a fixed seed $\sigma$ and the model’s nominal configuration. We refer to this value as the Fixed-Seed Sampling Likelihood (FSSL). This is simply the fixed-seed posterior distribution evaluated at a specific sampled token. Intuitively, this measures how consistent a token is with the model’s expected non-deterministic behavior under fixed-seed execution.
@@ 4.5 Solution 3: Classification using the Fixed-Seed Sampling Likelihood Policy
with $\tau$ calibrated on benign traffic to meet a target false–positive rate. Per-token decisions (or summed log-likelihoods) can be aggregated over sequences, users, or time windows for higher-level alerts. We instantiate estimates of the FSSLfor two samplers: we introduce Token-DiFR to estimate FSSLquantity for samplers that use the Gumbel-Max Trick, FSSL-GM, and we introduce Token-IPT-DiFR (FSSL-IPT) for samplers that use Inverse Probability Transforms for sampling, which estiamtes FSSL-IPT. Subsequent sections derive their estimators and evaluate performance.
@@ 4.6 Theoretical Exfiltration Limitations
Our proposed architecture allows us to provably bound the ability of any covert adversary to exfiltrate data out of the server under mild trust and cryptographic assumptions (see Appendix D). Recall that our dichotomy is to either have the adversary get caught or else limit the exfiltration rate of the adversary. This section is concerned with proving limits in the latter case. In other words, we will bound how quickly an adversary can exfiltrate without getting flagged. This is the best we can hope to prove in our context, without assumptions on the adversaries exfiltration pattern or frequency. We discuss later in this section how our theorems are related to the Exfiltration game outlined in early sections.
We prove three different bounds for the FSSL policy (Section 4.5), which apply under different architectures and assumptions. The first is the most general and applies to any adversary under the minimal assumptions explicated in Appendix D and using the simple FSSL filter we propose in Section 4.5. The remaining two are tighter bounds but only apply under additional assumptions, which can be enforced through particular architectures. We go on to discuss the plausibility and tightness of the provable bounds.
@@
First, we recall Theorem 2.2, which shows that any steganography scheme can only transmit as much information as is in the entropy of its output distribution. We combine this with two other simple observations to give rate limitation for our FSSL policy. Because we bound with resepct to such a simple policy, this bound will also apply to any the optimizations and augmented policies discussed in Section 5.
Theorem 4.1 is a most general rate-limiting bound. It applies to exfiltration adversaries where both the encoder and decoder have access to the seed that is used for generation. In the worst case, this bound shows that an adversary can only exfiltrate as many bits as the FSSL policy is set to. However, as we show in Section 6, for many prompts and seeds this bound is much tighter than the worst case, since $\mathsf{Supp}(\mathcal{D}_{\leq r})\ll 2^{r}$. Here, we use $\mathcal{D}_{\leq r}$ to indicate the honest distribution, restricted to the responses which have log-likelihood at most $r$. We also use $\mathsf{Supp}$ to indicate the set of responses which support the input distribution.
@@
For a given prompt and seed, let $\mathcal{D}$ be the honest response distribution, and let $\mathcal{A}$ be an adversary that produces a response distribution $\mathcal{D}_{\mathcal{A}}$ and exfiltrates $r^{\prime}$ bits per token on average. Then, $\mathcal{A}$ will be caught by the FSSL policy set at rate $r$ with probability $p:=\Pr_{t\sim\mathcal{D}_{\mathcal{A}}}[t\not\in\mathsf{Supp}(\mathcal{D}_{\leq r})]$, and when $\mathcal{A}$ is not caught $r^{\prime}\leq\log|\mathsf{Supp}(\mathcal{D}_{\leq r})|\leq r$.
@@
The FSSL policy flags any response which is too surprising. Formally, it computes the log-likelihood (a.k.a. surprisal or empirical entropy) for the given response and seed. Then, it tests the likelihood $p$ and the length of the response $\ell$ to check if $p/\ell\geq r$. Among the honest distribution, the only valid responses which are below this threshold are exactly $\mathsf{Supp}(\mathcal{D}_{\leq r})$. If the adversary has any response outside this set, then the policy will catch $\mathcal{A}$. Therefore, the probability they are caught probability exactly $\Pr_{t\sim\mathcal{D}_{\mathcal{A}}}[t\not\in\mathsf{Supp}(\mathcal{D}_{\leq r})]$.
Whenever $\mathcal{A}$ passes by the policy without being flagged then, it is because they output a response from the set $\mathsf{Supp}(\mathcal{D}_{\leq r})$. And, the maximum entropy a distribution can achieve on such a distribution is $\log|\mathsf{Supp}(\mathcal{D}_{\leq r})|$, and therefore the amount of information is also bounded (Theorem 2.2). Further, there can only be at most $2^{r}$ responses what appear with probability at least $2^{-r}$. So, in the worst case $r^{\prime}=r$. ∎
@@
Under some additional assumptions, one may expect that the decoding adversary cannot obtain the seed that used by the encoder. In this case, Theorem 4.1 is too loose. We can obtain stronger bounds if we make some additional assumptions about the surrounding infrastructure. We show two bounds under these additional assumptions below. See Appendix D for more discussion about these assumptions.
The following bounds can be applied for any policy that that has a notion of permitted tokens. These are the tokens that would be allowed by the policy for the given prompt, seed, and prefix. For example, the FSSL policy may allow any token which is not too surprising for the given seed, which is usually the 1 or 2 adjacent logits next to the one the verification server produced.
First, using the capacity limitations covered in Theorem 2.4, we obtain the bound below. However, this requires some other careful assumptions about the security and implementation of the seed expansion. In particular, we need that the server can’t expand the seed on their own and that they must commit to a token in order to get the randomness used for the next logit. This is the most interactive infrastructure we propose, but we can obtain a very strong bound.
@@
For a given prompt and policy, let $p$ be the probability over a random seed that token generation has $q>1$ permitted tokens. Then, any efficient adversary $\mathcal{A}$ that is not caught by the policy exfiltrates $r^{\prime}\leq\log q-h(1-p)-(1-p)\log(q-1)$ bits per token on average, where $h$ is the binary entropy function, assuming:
@@
    the seed is expanded adaptively as the response is generated,
@@
    the seed is longer than the response,
@@
    the seed is expanded by a cryptographic hash function, [^5] Formally, our bounds need the stronger assumption that we can model the hash function as a random oracle [CCS:BelRog93].
  4. 4.
    and the adversary $\mathcal{A}$ only receives one hash evaluation after committing to each token.
@@
Since the adversary is only given the random choice for the current token and must commit to the token before getting the next random choice, each logit is uniformly random and unpredictable by the sender. And, by assuming the seed is longer than the response, each of these choices remains uniformly random conditioned on the previous ones, since there isn’t enough response length to learn the seed.
Therefore, we can model each token as an independent channel and apply Theorem 2.4 directly. The sender must transmit $\ell$ bits to the receiver with no error. And, we consider each token sent as a channel, which either has $q>1$ permitted tokens (with probability $p$) or has one permitted token, determined by the seed and context (i.e., if the adversary were to deviate they would get caught). Theorem 2.4 gives a direct bound on the number of messages $n$ in this case. Specifically that $n\geq\ell/(\log q-h(1-p)-(1-p)\log(q-1)),$ where we model the probability that there is only one permitted token (1p)(1-p) as uniformly random symbol, since the sender has no control over it (recall that it is unpredictable). As we’re interested in the bits per token, we can therefore bound $\ell/n\leq\log q-h(1-p)-(1-p)\log(q-1)$ as claimed. ∎
If the interactivity of the hash function evaluation is too expensive, we prove one final, slightly weaker bound. In this case we need only assume that the adversary is computationally bounded. In practice, this assumption can be enforced by requiring the server output tokens at some minimum rate. This way the adversary has a limited amount of time to search for “good” responses.
Theorem 4.3.
For a given prompt and policy, let $p$ be the probability over a random seed that token generation has $q>1$ permitted tokens. Fix any constant $\varepsilon>0$. Then, any efficient adversary $\mathcal{A}$ that is not caught by the policy exfiltrates $r^{\prime}\leq(p+\varepsilon)\log q$ bits per token with high probability on each message, where $h$ is the binary entropy function, assuming:
  1. 1.
    the seed is expanded adaptively as the response is generated,
  2. 2.
    the seed is expanded by a cryptographic hash function with a random key unknown to $\mathcal{A}$,
  3. 3.
    the adversary $\mathcal{A}$ queries the hash function $O(k)$ times, where $k$ is the response length.
Proof.
We basically just need to show that at most $p+\varepsilon$ of the logits have more than $1$ permitted token. This is equivalent to bounding the probability of the adversary finding a response sequence where $(p+\varepsilon)\cdot k$ tokens have more than 1 permitted token. We can bound this using Lemma 4.4. Intuitively, each response has a probability $p$ of having more than $1$ permitted token (which we will model with $H$ outputting 0). Therefore, given $O(k)$ queries, the adversary can only find a sequence like this with probability at most $e^{-O(k)}$ which is negligible in the response length. So, with high probability, the adversary has only $(p+\varepsilon)k$ locations to exfiltrate on, each with an alphabet of size $q$. So, the maximum number of bits that can be exfiltrated is $(p+\varepsilon)\log q$ per token. ∎
@@
Let $\mathcal{A}$ by an algorithm issuing at most $N=O(k)$ queries to a random function $H:\{0,1\}^{*}\to\{0,1\}$ that is 0 with probability $p$, which outputs a string $x\in\{0,1\}^{k}$. Then, the probability that $p^{\prime}k$ of the prefixes of $x$ evaluate to 0 is at most $e^{-\Theta(k)}$.
@@
Here we provide some estimates on how tight we expect the above bounds to be against a motivated adversary. Theorem 4.1 is believed to be tight in theory. As long as basic cryptography exists, steganography exists as a consequence of Theorem 2.3 [C:HLvA02]. Even in practice, an adversary that carefully designs a steganographic scheme against the FSSL policy should likely be able to get close to this bound. They could likely draw heavily on techniques from steganographic literature to design schemes.
Theorems 4.2 and 4.3 are likely considerably looser under the stated assumptions. The bound allows an adversary to choose from any of the permitted tokens. So an adversary would likely need to consistently choose rare tokens to try to meet the stated limit. But even for a simple policy, like the FSSL policy, an adversary would need to be much more careful about their token choice across the response to avoid getting flagged. Moreover, it’s also not clear how to design a steganographic scheme for this setting. When the seed is known, both parties have access to the distribution, the typical setting for steganography. But, when the seed is secret, one may need to design a model and policy specific encoding scheme to try and achieve a meaningful rate.
@@
Relationship to the Exfiltration Game.
Finally, we sketch how use our theorems to prove bounds for any adversary’s success in Experiment 3.2. We can use a generic transformation from per-response rate limits to limits in the experiment. For this bound, we specifically focus on the FSSL policy and proving that the policy wins the Exfiltration game ($\mathsf{ExfilGame}$) with good probability under some reasonable plausibility assumptions.
Notice that all three bounds we have the following form: a response $y_{t}$ that does not raise an alarm has capacity $C_{t}$ except with probability at most $p_{t}$. (These values may depend on the response-length of $y_{t}$). So, where inspections are randomly selected, any adversary that exfiltrates more than $C>\log(M)+\sum_{t}C_{t}$ bits must have some query $t^{\prime}$ on which they exfiltrated more than $C_{t}$. [^6] Notice that this gives the adversary $\log(M)$ bits “for-free” by allowing them to adaptively choose whether a response is honest or dishonest. It is unclear whether this capacity could be exploited in practice, but for our bounds we give the adversary this power.
More carefully, let $E_{1}$ be the event that $\mathcal{A}$ correctly guesses $\mathbf{m}$, $E_{2}$ be the event that $\mathcal{W}$ raises the alarm on a dishonest message, and $E_{3}$ be the event that the total capacity of $\mathcal{A}$’s responses exceeds $C$.
$\Pr[\mathsf{ExfilGame}_{\mathcal{A},\mathcal{W}}(n,M,B_{I},B_{A})=1]$ $=\Pr[E_{1}\land\neg E_{2}]$
$=\Pr[E_{1}\land\neg E_{2}\land\neg E_{3}]+\Pr[E_{1}\land\neg E_{2}\land E_{3}]$
$\leq\Pr[E_{1}|\neg E_{3}]+\Pr[\neg E_{2}|E_{3}]$
$=\Pr[E_{1}|\neg E_{3}]+(1-\Pr[E_{2}|E_{3}])$
Then, we take $p_{t}=0$ for simplicity (and because this is true for Theorem 4.1), so that the warden will always raise the alarm if message $t$ exceeds $C_{t}$. Now, because we know that when $E_{3}$ occurs, there is some $t^{\prime}$ which exceeds the exfiltration rate, then $\Pr[E_{2}|E_{3}]$ is at least $B_{I}$ minus the probability that all $B_{A}$ alarms are raised before $t^{\prime}$ is encountered. Then using the Chernoff-Hoeffding bound, this is at most $\exp\big(-B_{I}D_{\mathrm{KL}}\big(\frac{B_{A}}{B_{I}}\|\alpha\big)\big)$, where $\alpha<B_{A}/B_{I}$ is the false positive rate of $\mathcal{W}$ on the honest distribution. We can also use the well-known information theoretic fact that a capacity $C$ bounded channel has limited ability to predict a random variable, which gives $\Pr[E_{1}|\neg E_{3}]\leq 2^{C-n}$. Together this gives us
$\Pr[\mathsf{ExfilGame}_{\mathcal{A},\mathcal{W}}(n,M,B_{I},B_{A})=1]\leq\frac{M-B_{I}}{M}+2^{C-n}+\exp\big(-B_{I}D_{\mathrm{KL}}\big(\frac{B_{A}}{B_{I}}\|\alpha\big)\big),$
which is near optimal asymptotically, since $\mathcal{A}$ can always choose a maximal exfiltration limit on a few messages and hope that $\mathcal{W}$ does not inspect those messages. Note also that in this argument, we’ve assumed that each response is independent, which requires some assumption on the seeds and cryptography that is used.

### v3

@@ 5 Theoretical Framework: Verification System
Figure 4: Overview of the verification system for runtime output validation. A lock symbol denotes parts of the inference verification scheme that require high integrity, further explored in Section E
@@
    Inference Logging. The API Server logs all inference inputs and outputs, together with references to the exact model release, code version, and the seed used for that inference. [^5] In the event that not every (input, output, seed) tuple is verified, randomness is critical in the process of selecting which messages to verify is essential. If the messages-to-be-verified is predictable, an adversary could selectively encode sensitive data in unverified outputs.
@@
The degree of isolation of the Verification Server is a tunable security parameter. In minimal deployments, this can be enforced through access controls and containerization; in the extreme, the verification server may be fully air-gapped. All trust assumptions are detailed in Appendix E.
5.1 Sampled-Token Verification via Exact Matching
@@
For efficiency, the same idea can be executed without iterative decoding. Using prefill, a single forward pass over the input+output yields the next-token distributions at every position; sampling (with the recorded seed) and exact matching can then be done position-wise in parallel. This preserves teacher-forcing semantics while running significantly faster. Full pseudocode appears in Appendix C.3.
5.2 Classification using the Fixed-Seed Sampling Likelihood Policy
In Section 3.4, we introduced the fixed-seed posterior distribution, which captures the stochastic variability of model outputs that can arise under honest non-determinism. Building on this, we define a practical decision rule for distinguishing between valid and suspicious generations. Our approach centers on estimating the likelihood that a given output token $t$ was honestly generated under a fixed seed $\sigma$ and the model’s nominal configuration. We refer to this value as the Fixed-Seed Sampling Likelihood (FSSL). This is simply the fixed-seed posterior distribution evaluated at a specific sampled token. Intuitively, this measures how consistent a token is with the model’s expected non-deterministic behavior under fixed-seed execution.
@@ 5.2 Classification using the Fixed-Seed Sampling Likelihood Policy
with $\tau$ calibrated on benign traffic to meet a target false–positive rate. Per-token decisions (or summed log-likelihoods) can be aggregated over sequences, users, or time windows for higher-level alerts. We instantiate estimates of the FSSLfor two samplers: we introduce Token-DiFR to estimate FSSLquantity for samplers that use the Gumbel-Max Trick, FSSL-GM, and we introduce Token-IPT-DiFR (FSSL-IPT) for samplers that use Inverse Probability Transforms for sampling, which estimates FSSL-IPT. Subsequent sections derive their estimators and evaluate performance.
@@ 5.3 Theoretical Exfiltration Limitations
Our proposed architecture allows us to provably bound the ability of any covert adversary to exfiltrate data out of the server under mild trust and cryptographic assumptions (see Appendix E). Recall that our dichotomy is to either have the adversary get caught or else limit the exfiltration rate of the adversary. This section is concerned with proving limits in the latter case. In other words, we will bound how quickly an adversary can exfiltrate without getting flagged. This is the best we can hope to prove in our context, without assumptions on the adversaries exfiltration pattern or frequency. We discuss later in this section how our theorems are related to the Exfiltration game outlined in early sections.
We prove three different bounds for the FSSL policy (Section 5.2), which apply under different architectures and assumptions. The first is the most general and applies to any adversary under the minimal assumptions explicated in Appendix E and using the simple FSSL filter we propose in Section 5.2. The remaining two are tighter bounds which require additional assumptions.We go on to discuss the plausibility and tightness of the bounds.
@@
First, we recall Theorem B.2, which shows that any steganography scheme can only transmit as much information as is in the entropy of its output distribution. We combine this with two other simple observations to give rate limitation for our FSSL policy. Because we bound with respect to such a simple policy, this bound will also apply to any the optimizations and augmented policies discussed in Section 6.
Theorem 5.1 is a most general rate-limiting bound. It applies to exfiltration adversaries where both the encoder and decoder have access to the seed that is used for generation. In the worst case, this bound shows that an adversary can only exfiltrate as many bits as the FSSL policy is set to. However, as we show in Section 7, for many prompts and seeds both the false positive rate and exfiltration limitations are much tighter than the stated in this worst case, since $\mathsf{Supp}(\mathcal{D}_{\leq\tau})\ll 2^{r}$. Here, we use $\mathcal{D}_{\leq\tau}$ to indicate the honest distribution, restricted to the responses which have log-likelihood at most $r$. Here, we’re using $\mathsf{Supp}$ to indicate the set of responses which support the input distribution.
@@
For any distribution $\mathcal{D}$, $\mathcal{W}_{\mathrm{FSSL},\tau}$ at threshold $\tau$ is $(\alpha,\varepsilon)$-exfiltration secure for $\alpha=|\mathcal{T}|^{\ell}\cdot 2^{-\ell\cdot\tau}$ and $\varepsilon=(|\mathsf{Supp}(\mathcal{D}_{\leq\tau})|-1)/2^{n}\leq(2^{\ell\cdot\tau}-1)/2^{n}$.
@@
The $FSSL$ policy flags any response which is too surprising. Formally, it computes the log-likelihood (a.k.a. surprisal or empirical entropy) for the given response and seed. Then, it tests the likelihood $p$ and the length of the response $\ell$ to check if $p/\ell\geq\tau$. Among the honest distribution, the only honest responses which are below this threshold are exactly $\mathcal{S}:=\mathsf{Supp}(\mathcal{D}_{\leq r})$.
First, notice that by definition, for any $\hat{T}\not\in\mathcal{S}$, $\Pr_{T\sim\mathcal{D}}[T=\hat{T}]<2^{-\ell\cdot\tau}$ and therefore in total, we have that $\alpha\leq|\mathcal{T}|^{\ell}\cdot 2^{-\ell\cdot\tau}$ which establishes our claim about $\alpha$.
Next, observe that whenever the adversary has any response outside $\mathcal{S}$, then the $\mathbf{Exfil}$ game outputs $\bot$. Therefore, when $\mathcal{A}$ is not flagged by the policy, it is because their message $\hat{T}$ is in the set $\mathcal{S}$. Notice that the maximum entropy a distribution can achieve on such a distribution is $\log|\mathcal{S}|$, and therefore the capacity of a channel with such a support size is also bounded by $|\mathcal{S}|$ (Theorem B.2). The final inequality comes from the observation that there are at most $2^{\ell\cdot\tau}$ responses what appear with probability at least $2^{-\ell\cdot\tau}$, which establishes the $\varepsilon$ security claim. ∎
@@
Under additional assumptions or infrastructure, one may expect that the decoding adversary cannot obtain the seed that used by the encoder. In this case, Theorem 5.1 is too loose, and we should aim to achieve bounds in the restricted game $\mathbf{Exfil}^{\mathbf{r}}$ (Experiment 3). In this section, we show two bounds in this restricted game below. See Appendix E for more discussion about these assumptions.
The bounds in this section attempt to generalize beyond the $FSSL$ policy and will apply to any policy with a particular structure. Motivated by our experimental results, we see that each token of a generation is nearly deterministic given the prompt and seed. In particular only about $2\%$ of tokens in our experiments are “branching points,” where floating-point noise makes the next token undetermined.
Permitted tokens.
The above observation motivates our characterization of policies $\mathcal{W}$ where at each index $i$, there is only one token $t$ which is permitted, i.e., if $\hat{T}[i]\neq t$ then $\mathcal{W}(\hat{T})=1$. In practice, we don’t quite obtain full determinism, so we instead define $(p,q)$-permitted token policies. Where, at each index there is one permitted token with probability (1p)(1-p), or else there are $q>1$ permitted tokens with probability $p$ (and these probabilities are independent).
Based on our experiments, we can approximate the exfiltration bounds of the $FSSL$ policy with $p=0.02$ and $q=2$. However, we note that this is a very simplified abstraction for the purposes of obtaining a possible stronger bound under additional assumptions. In the real world, it is unclear how closely policies satisfy the $(p,q)$-definition. It is possible an adversary could correlate where permitted tokens occur (breaking independence) or choose an adversarial prompt which has additional “branching points,” which may be interesting for future work to explore.
First, using the capacity limitations covered in Theorem B.4, we obtain the bound below. However, this requires some other careful assumptions about the security and implementation of the seed expansion. In particular, we need that the adversary can’t expand the seed on their own and that they must commit to a token in order to get the randomness used for the next logit. This is the most interactive infrastructure we propose, but we can obtain a very strong bound.
@@
Let $\mathcal{W}$ be a $(p,q)$-restricted policy with false positive rate $\alpha$ on $\mathcal{D}$. Define $r^{\prime}:=\log q-h(1-p)-(1-p)\log(q-1)$ where $h$ is the binary entropy function. Then, $\mathcal{W}$ is $(\alpha,\varepsilon)$-secure against admissible adversaries for $\varepsilon=(2^{\ell\cdot r^{\prime}}-1)/2^{n}$, under the following additional conditions:
@@
    the seed is longer than the response, i.e., $\lambda>\ell\log|\mathcal{T}|$,
@@
    and the adversary $\mathcal{A}_{1}$ computes $\mathcal{D}^{\mathsf{H}}(\sigma)$ once and never queries $\mathsf{H}$ otherwise.
Proof.
Since the adversary $\mathcal{A}_{1}$ is only given the honest $\mathsf{H}$ evaluation while streaming an honest response, each logit is uniformly random and unpredictable by $\mathcal{A}_{1}$. And, by assuming the seed is longer than the response ($\lambda>\ell\log|\mathcal{T}|$), each of these choices remains uniformly random conditioned on the previous ones, since there isn’t enough response length to learn the seed.
Therefore, we can model each token as an independent channel and apply Theorem B.4 directly. Adversary $\mathcal{A}_{1}$ must transmit $n$ bits to the receiver with no error. And, we consider each token sent as a channel, which either has $q>1$ permitted tokens (with probability $p$) or has one permitted token, determined by the seed and context (i.e., if the adversary were to deviate they would get caught). Theorem B.4 gives a direct bound on the number of messages $M$ in this case. Specifically, we have $M\geq n/(\log q-h(1-p)-(1-p)\log(q-1)),$ where we model the probability that there is only one permitted token (1p)(1-p) as uniformly random symbol, since the sender has no control over it (recall that it is unpredictable). As we’re interested in the bits per token, we can therefore bound $n/M\leq\log q-h(1-p)-(1-p)\log(q-1)$ as claimed. ∎
If the interactivity of the hash function evaluation is too expensive, we prove one final, slightly weaker, bound. In this case we need only assume that the adversary is computationally bounded. In practice, this assumption can be enforced by requiring the server output tokens at some minimum rate. This way the adversary has a limited amount of time to search for “good” responses.
Theorem 5.3.
Let $\mathcal{W}$ be a $(p,q)$-restricted policy with false positive rate $\alpha$ on $\mathcal{D}$. For any constant $\delta>0$ define $r^{\prime}:=(p+\delta)\log q$. Then, $\mathcal{W}$ is $(\alpha,\varepsilon)$-secure against admissible adversaries for $\varepsilon=(2^{\ell\cdot r^{\prime}}-1)/2^{n}+e^{(-O(\ell))}$, under the following additional conditions:
  1. 1.
    the seed is longer than the response, i.e., $\lambda>\ell\log|\mathcal{T}|$,
  2. 2.
    The inference distribution $\mathcal{D}^{\mathsf{H}}(\sigma)$ samples randomness sequentially,
@@
    and the adversary $\mathcal{A}_{1}$ queries $\mathsf{H}$ at most $O_{\delta}(\ell)$ times.
@@
The main goal of this result amounts to showing that at most $p+\delta$ fraction of the logits in $\mathcal{A}_{1}$’s output $\hat{T}$ have more than one permitted token. This is equivalent to bounding the probability of the adversary finding a response sequence where $(p+\delta)\cdot\ell$ tokens have more than 1 permitted token. We can bound this using Lemma 5.4 below. Intuitively, each response has a probability $p$ of having more than $1$ permitted token (which we will model with $\mathsf{H}$ outputting 0). Therefore, given $O(\ell)$ queries, the adversary can only find a sequence like this with probability at most $e^{-O(\ell)}$ which is negligible in the response length. So, with high probability, the adversary has only $(p+\delta)\ell$ locations which are not deterministic, each with an alphabet of size $q$. So, the maximum number of bits that can be exfiltrated is bounded by $\ell(p+\delta)\log q$. ∎
@@
Let $\mathcal{A}^{\mathsf{H}}$ by an algorithm issuing at most $N=O(k)$ queries to a random function $\mathsf{H}:\{0,1\}^{*}\to\{0,1\}$ that is 0 with probability $p$, which outputs a string $x\in\{0,1\}^{k}$. Then, the probability that $p^{\prime}k$ of the prefixes of $x$ evaluate to 0 is at most $e^{-\Theta(k)}$.
@@
Here we provide some estimates on how tight we expect the above bounds to be against a motivated adversary. Theorem 5.1’s exfiltration bound $\varepsilon$ is likely to be tight in theory. As long as basic cryptography exists, steganography exists as a consequence of Theorem B.3 [HLV02]. Even in practice, an adversary that carefully designs a steganographic scheme against the FSSL policy should likely be able to get close to this bound. They could likely draw heavily on techniques from steganographic literature to design schemes.
Theorems 5.2 and 5.3 are likely considerably looser under the stated assumptions. The bound allows an adversary to choose from any of the permitted tokens. So an adversary would likely need to consistently choose rare tokens to try to meet the stated limit. But even for a simple policy, like the FSSL policy, an adversary would need to be much more careful about their token choice across the response to avoid getting flagged. Moreover, it’s also not clear how to design a steganographic scheme for this setting. When the seed is known, both parties have access to the distribution, the typical setting for steganography. But, when the seed is secret, one may need to design a model and policy specific encoding scheme to try and achieve a meaningful rate.